# CAAP: Context-Aware Action Planning Prompting to Solve Computer Tasks with Front-End UI Only

### v1

@@ 4 CAAP prompting
The performance of LLM-based agents is heavily dependent on how the prompts are structured. By focusing on delivering relevant contextual information in the most efficient manner and inducing CoT reasoning during response generation, our prompting technique significantly enhances the decision-making capabilities of LLM agents in computer tasks. CAAP prompts, as summarized in Figure 3, consist of four main parts.
Human demonstrations Exemplars of action trajectory with rationale
Surrounding context information Task description The last LLM response Action trajectory Visual observation description Candidates of action types
CoT-inducing instructions Instruction for reviewing action trajectory Instruction for reviewing visual observation Instruction for improving action plan
Instruction of decision making for next actions Guidance for the response format of actions Request for decision making of the following actions Extra guidelines
Figure 3: Content design for the CAAP prompting. Figure 9 and 10 in Appendix F show an example of a prompt generated during the MiniWob++ experiments and the LLM’s response to this prompt.
@@
It is widely recognized that providing few-shot exemplars significantly enhances the advanced reasoning capabilities of LLMs. The CAAP prompt begins with a list of actions performed by a human demonstrator solving a task similar to the given objective, transcribed in plain text. In practical scenarios, the Retrieval-Augmented Generation (RAG) method can be employed to extract similar task records from a database. For MiniWoB++ experiments, we use tasks of the same type, created using different seeds. The effectiveness of the demonstration text can be greatly improved by incorporating rationale texts behind each action, even when those rationales are not from the demonstrators themselves but are autonomously generated by an LLM (see Appendix A), as demonstrated in Figure 5.
@@
This second part of the CAAP prompt constitutes the core, including essential elements crucial for the language model to make accurate decisions. It encompasses a description of the target task, state descriptions from the Visual Observer, a record of past actions, and the available range of actions for the agent. For continuity in the decision-making process, this section also includes a response from the agent that recaps its previous action plan, ensuring consistency in the strategy used to pursue the goal.
@@
The CoT-inducing instructions section contains phrases that encourage the LLM agent to interpret the surrounding context information independently, thereby inducing the Chain of Thought. Instruction for reviewing action trajectory encourages the agent to review its actions taken so far, enhancing its understanding of the connection between these past actions and the current state of the environment. Instruction for reviewing visual observation prompts the agent to examine the current state’s observations, directing its focus on relevant objects needed to complete the task. Instruction for improving action plan requests the development of a follow-up action plan, considering the self-generated texts so far.
Instruction of decision making for next actions
The final section of the CAAP prompt primarily features phrases designed to safeguard the LLM’s output while requesting the next action. It incorporates formatting guidelines and additional directives specifically tailored to reduce common errors made by the particular LLM in use.
@@ 5 Experiments
5.1 Experiment setup
We evaluate the performance of our CAAP-based agent within the MiniWoB++ environment (Liu et al. (2018); Min ). MiniWoB++ is designed with over 100 web interaction task scenarios that challenge software agents to formulate and execute effective action sequences. These scenarios encompass a variety of UI interactions, including clicking buttons, filling out forms, and performing drag-and-drop tasks. Although MiniWoB++ offers numerous APIs that could streamline agent-environment interactions, we deliberately avoid using these artificial system shortcuts, which are typically absent in everyday desktop tasks.
@@ 5.1 Experiment setup
The list of tasks supported by MiniWoB++ has been continuously expanding. Previous studies using this benchmark often choose subsets of the tasks that are most suitable for demonstrating their methods. For our experiment, we have selected 67 tasks based on the following criteria: (1) the types of UI elements used in the task are known to our Visual Observer module, as listed in Table 2 of Appendix B (2) the UI elements appearing while solving the task do not extend beyond the default screen area of 160x210 pixels (3) the task does not require color recognition, as we have yet to train our Visual Observer module for this capability .
For model training [^2] The UI element detection and understanding models are fine-tuned with this dataset. While the former utilizes Ultralytics YOLOv8.0.143 (ult ) with the default hyper-parameters except ’box’ of 0.9, ’cls’ of 0.05, and ’dfl’ of 0.05, the latter employs the base model downloaded from https://huggingface.co/google/pix2struct-base without any configuration changes., we use two types of datasets. The first dataset is for fine-tuning the models in Visual Observer. To prepare this dataset, two individuals spent 10 hours each, collecting the screenshots and annotating the UI elements within. The second dataset is the record of human demonstrations solving MiniWob++ tasks. This demo data was collected in 2 hours by one individual. Detail on datasets and the model training is described in detail in Appendix A.
To differentiate between the train datasets and the test dataset, we assign distinct seed numbers when generating individual task instances within MiniWoB++. Image annotation data was collected from tasks with seed numbers ranging from 1000 to 2999, while demonstration data was sourced from seeds ranging from 3000 to 3999. Testing was conducted using seeds ranging from 0 to 999. For testing, we conducted at least 50 repetitions for each experiment to guarantee the statistical reliability.
@@
For the decision-making language model, we utilize the Azure OpenAI gpt-4-0125 model. This model features the ‘function calling’ capability, thus our prompt does not directly include the list of action types (as depicted in Figure 3). Instead, this contextual information is conveyed to the language model through arguments in the function calling feature. An example of how the function calling is applied is illustrated in Figure 11 in Appendix G.
5.2 Experiment result
Table 1 and Figure 4 summarize the performance of our agent in comparison with other methods documented in the literature for solving MiniWoB++ tasks. Of the 104 tasks solved by CC-Net (Humphreys et al. (2022)), our comparison focuses on the 100 tasks available in the latest version of MiniWoB++. The success rates listed in Table 1 are calculated based on the task coverage of each method. Reducing task coverage can artificially inflate success rates, as it allows the inclusion of only those tasks with high success rates. Consequently, We introduces a new metric in Figure 4 for comparing research outcomes in a fairer way, which involves counting the number of tasks with success rates exceeding 80% and 90%. Additionally, we employ a second metric that recalculates success rates by assuming all attempts have failed for tasks that are unsupported or unreported.
Table 1: Comparison with other methods that solve MiniWoB++ problems. FT, DM, VM stand for Fine-Tuning, Decision Model, Vision Model, respectively. Task success rate for each task is shown in Table 6 in Appendix H.2
Method Modality Model Training Datasize Tasks Success
Vision (VM) Decision (DM) Image Demo Rate
RCI HTML - gpt-4 - - < 0.1K 54 94.0%
AdaPlanner HTML - text-davinci-003 - - < 0.1K 53 92.9%
WebGUM HTML+Image ViT-B16 Flan-T5-XL FT (both) - 401K 57 94.2%
CCNet DOM+Image ResNet Transformer SL+RL (DM) - 2.4M 100 93.6%
Image < 25.0%
Pix2Act Image Pix2Struct Transformer SL+RL (DM) - 1.3M 58 96.1%
SeeClick Image ViT Qwen-VL 9.6B FT (DM) - 2.8K 55 69.4%
CAAP
Image
YOLOv8
+Pix2Struct
gpt-4-0125 FT (VM) 1.7K < 0.1K 67 94.4%
Human Image - - - - - 100 93.5%
Figure 4: Comparison on the number of stably solved MiniWoB++ tasks and on the overall success rate. The success rates on the red line are calculated by assuming uncovered tasks have all failed among the total coverage of all methods. The precise values are described in Table 5 of Appendix H.1.
One of the most notable methods compared is CC-Net, which addresses nearly all tasks in the MiniWoB++ benchmark. However, its development required an extensive labor input of 6,300 hours to collect 2.4M demos, and its high dependency on DOM information poses challenges for practical application. RCI (Kim et al. (2023)) and AdaPlanner (Sun et al. (2023)), while reporting more than 92% task-solving success rates with fewer than 100 demos, rely heavily on HTML input, which limits their task coverage. When recalculating success rates including unsupported or unreported tasks, their success rates drop to 50.8% and 49.2%, respectively. WebGUM (Furuta et al. (2023)) combines HTML and image inputs, which has improved task-solving success rates by eliminating constraints on task coverage. Despite utilizing a dataset of 401K demos, its results are only slightly better compared to RCI. Pix2Act (Shaw et al. (2023)), the first method to achieve high success rates in MiniWoB++ tasks using only image inputs, covers 58 tasks with over 80% success in 55 types, supported by a large dataset of 1.3M demos. Conversely, SeeClick, which also utilizes only image inputs, reports SL results with a significantly smaller demo dataset, yet its success rate falls to 38.2% for the 100-types metric.
Our method ranks as the most advanced among image-only agents, outperforming Pix2Act by covering 67 tasks, with 61 of these tasks achieving success rates above 80%. Additionally, it secures a 63.3% success rate, which is 7.5% higher than Pix2Act’s 55.8%. This substantial improvement, achieved with fewer than 100 demonstrations, is attributed to the powerful reasoning capabilities of an LLM and our meticulously structured agent design.
Ablation studies
Figure 5 presents experiment results comparing the effects of human demonstrations and CoT inducing instructions in our CAAP prompting method. It is shown that the use of CoT inducing instructions results in a meaningful improvement in performance. Removing CoT inducing instructions from the prompt results in a 2.4% drop compared to the baseline, and 4.8% drop when the human demo is absent. When only rationales are excluded from human demonstrations, there is a reduction in task-solving success rate by 3.3% relative to the fully-prompted baseline. Excluding the whole human demonstrations leads to a decrease of 8.7%. Additionally, the removal of both CoT-inducing instructions and human demonstrations results in an decrease in success rate by as much as 13.5%, demonstrating that effective ICL provided by our CAAP promping scheme significantly enhances the decision-making capacity of the LLM model in use.
Comparison with the RCI prompting
We conducted additional experiments to isolate the effect of CAAP prompting by comparing a standard RCI agent with another enhanced solely by our CAAP prompting mechanism. Both agents maintained the same input and action domains; only the prompting mechanism differed. We executed the original code from the RCI official GitHub without modifications, testing only tasks listed in the RCI GitHub repository. We ensured consistency in task instances by using identical seeds and employed the same demo set provided by the repository. The list of tasks and their respective success rates are documented in Table 7 of Appendix H.2. The results demonstrate that applying CAAP prompting to the RCI agent significantly increased the average task-solving success rate by 7.1%, from 76.9%333 An average task-solving accuracy of 76.9% was measured, which deviates from the results reported in the literature (Kim et al. (2023)). Despite diligent efforts to replicate the experimental setup exactly, the results documented in the referenced study could not be reproduced. Upon inquiry, the authors indicated that they no longer had access to the complete experimental logs that were used in their publication. to 84.0%. This experiment underscores the effectiveness of our CAAP prompting, confirming its potential to enhance the performance of LLMs in computing tasks.
@@ 6 Conclusion
In this paper, we introduce an agent designed to emulate human methods for solving computer tasks, specifically addressing the limitations of existing models reliant on HTML or DOM inputs, and those that combine supervised learning (SL) and reinforcement learning (RL), which are often complex and costly. Our agent operates solely on visual inputs and utilizes a large language model (LLM). We also introduce the CAAP prompting approach to enhance the decision-making capabilities of ICL-based agents. Our evaluations using the MiniWoB++ benchmark demonstrate the superiority of our method. We critique existing performance measurements on this benchmark as biased and propose a more equitable benchmark for comparison. Our findings indicate that our agent can solve a broader range of tasks with high success rates using significantly less data than most other methods.
Furthermore, our agent exemplifies how configurations designed to mimic human behavior with minimal interface requirements can be effectively utilized in more universal settings, including robotics and mobility environments that rely on visual information. Designed to operate independently of specific environments, our approach offers broader applicability and significant advantages from a generality perspective.
Limitations
While our agent has proven effective in the MiniWoB++ benchmark, the scope of its validation remains limited. For a more comprehensive assessment from a research perspective, it is essential to test the agent across a broader array of benchmarks. Future research will aim to evaluate our agent using diverse platforms such as WebShop (Yao et al. (2022b)), AlfWorld (Shridhar et al. (2020)), and WebArena (Zhou et al. (2023)).
@@
Appendix A Train datasets

### v2

@@ 4 CAAP prompting
Human demonstrations Exemplars of action trajectory with rationale Surrounding context information Task description Action trajectory Visual observation description Candidates of action types CoT-inducing instructions Instruction for reviewing action trajectory Instruction for reviewing visual observation Instruction for improving action plan Instruction for deciding next actions Extra guidelines
Figure 3: Content design for the CAAP prompting. Figure 8 and 9 in Appendix H show an example of a prompt generated during the MiniWob++ experiments and the LLM’s response to this prompt.
The performance of LLM-based agents is heavily dependent on how the prompts are structured. By focusing on delivering relevant contextual information in the most efficient manner and inducing CoT reasoning during response generation, our prompting technique significantly enhances the decision-making capabilities of LLM agents in computer tasks. CAAP prompting constructs the prompt using the content depicted in Figure 3.
@@
Providing few-shot exemplars within the prompt to an LLM is a common practice that enhances the advanced reasoning capabilities of LLMs. Our CAAP prompt begins with few-shot examples, which are sequences of actions performed by human demonstrators solving tasks similar to the given objective. In practical scenarios, the Retrieval-Augmented Generation (RAG) method can be employed to extract similar task records from a database. The effectiveness of this demonstration portion of the prompt can be significantly enhanced by incorporating rationale texts behind each action, even when those rationales are not provided by the demonstrators themselves but are autonomously generated by an LLM, as shown in Section 5.3.
@@
This second part of the CAAP prompt forms the main section, containing all the essential information crucial for the language model to make accurate decisions. It includes a description of the target task, a record of past actions, state descriptions from the visual observer, and the available range of actions for the agent.
@@
The CoT-inducing instructions section contains phrases that encourage the LLM agent to interpret the surrounding context information independently, thereby inducing the Chain of Thought. Instruction for reviewing action trajectory encourages the agent to review its actions taken so far, enhancing its understanding of the connection between these past actions and the current state of the environment. Instruction for reviewing visual observation prompts the agent to examine the current state’s observations, directing its focus on relevant objects needed to complete the task. Instruction for improving action plan requests the development of a follow-up action plan, considering the self-generated texts so far. Lastly, the Instruction for deciding next actions provides guidance on the response format and requests a list of actions to be taken in the next steps.
Extra guidelines
The final section of the CAAP prompt primarily features phrases designed to safeguard the LLM’s output. It incorporates additional directives specifically tailored to reduce common errors made by the particular LLM in use.
@@ 5 Experiments
5.1 MiniWoB++
MiniWoB++ is an extension of the OpenAI MiniWoB benchmark, introduced in 2017, designed to evaluate an agent’s ability to handle a wide range of scenarios that may occur on a computer screen (Liu et al., 2018; Shi et al., 2017; Min, ). It consists of over a hundred of web interaction tasks that challenge the agent to devise and execute effective sequences of actions. These tasks include various UI interactions, such as button clicks, form filling, and drag-and-drop operations. We utilize the MiniWoB++ benchmark to thoroughly assess our agent’s effectiveness in navigating and solving diverse tasks within a simulated computer environment.
@@ 5.1 MiniWoB++
Previous studies using the MiniWoB++ benchmark often choose subsets of tasks most suitable for demonstrating their methods. We have also selected a total of 73 MiniWoB++ tasks that meet the following criteria: (1) tasks that consist of UI elements of types recognizable by our visual observer module, as listed in Table 7 of Appendix C (2) tasks where none of UI elements extend beyond the default screen area of 160x210 pixels (3) tasks that do not require color recognition, as our visual observer module is not yet trained for this capability.
For each selected task, we run at least 50 repetitions with different seeds to estimate the Success Rate (SR) for that task. Across all task types, we consistently use the identical CAAP prompt template without making any manual adjustments for specific tasks, unlike some other works (Kim et al., 2023).
Models and datasets
YOLO and Pix2Struct models within the visual observer have been fine-tuned to enable accurate recognition of images presented in MiniWoB++ environment. The image dataset for training has been collected through a process of capturing screenshots and annotating the UI elements within. This dataset consists of 1.76K screenshot images, gathered and annotated over a span of 10 hours by two individuals.
For human expert demonstrations, we generate between 0 and 5 demonstrations per task, tailored to the task’s complexity, amassing a total of 97 demonstrations for our experiments. We record the trajectory of each action performed by the demonstrators, alongside screenshots taken at the time of each action. These screenshots are subsequently analyzed to extract UI elements, enabling the identification of the UI component interacted with during mouse-related actions, such as clicks. This demonstration data have been collected by a single individual over a span of approximately two hours.
@@
For the decision-making language model, we utilize the Azure OpenAI gpt-4-0125 model. The Azure OpenAI API features the ‘function calling’ capability, thus our prompt does not directly include the list of action types (as depicted in Figure 3). Instead, this contextual information is conveyed to the language model through arguments in the function calling feature. An example of how the function calling is applied is illustrated in Figure 10 in Appendix I.
Experiment result
Table 1: Comparison with other image-only methods for solving MiniWoB++ problems.
Method Modality Image datasize Demo datasize Reported SR Reported tasks
Human Image - - 93.5% 104
CC-Net (no DOM) Image 0 2.4M 24.0% 104
Pix2Act Image 0 1.3M 96.2% 59
SeeClick Image 0 2.8K 69.4% 55
CAAP Image 1.8K 0.1K 94.5% 73
Table 1 compares the performance of our approach with other studies reported on the MiniWoB++ benchmark. Since each study reports results over different sets of tasks (See Figure 11 of Appendix K.), we have included both the SRs reported in the literature and the number of tasks used in each case. First, CC-Net, which covers the largest set of tasks, was trained on 2.4M human demonstration data across 104 tasks. However, when using only image input (excluding DOM), its performance dropped to 24%, indirectly highlighting the challenge of solving computer tasks using image inputs alone. Pix2Act, trained on 1.3M human demonstrations across 59 tasks, achieved an SR of 96.2%, demonstrating the feasibility of agents solving computer tasks with image input only. SeeClick, despite using a relatively small dataset of 2.8K demonstrations to learn click locations on screenshots, achieved an SR of only 69.4% across 55 tasks, showing limited effectiveness. In contrast, our approach, using only 97 demonstration samples, achieved a high task SR of 94.5% across 73 tasks. Although an additional 1.76K samples***These image samples were annotated via a partially automated process, making the data preparation process relatively efficient. were required to train the model for image input interpretation due to the modular nature of our approach, our method still demonstrated its effectiveness on a broader range of computer tasks with about 1.86K samples in total. Task-level SRs are detailed in Table 12 of Appendix J.3.
Of the 73 tasks included in our experiments, six tasks—drag-circle, drag-single-shape, find-greatest, generate-number, odd-or-even, and sign-agreement—were excluded from agent’s training, as they were not covered in the other studies. These tasks were deliberately chosen to test our agent’s ability to adapt to new, unseen tasks. Impressively, our agent successfully completed them with an average success rate (SR) of 95.3%. These results highlight the advantages of our modular architecture. Unlike other approaches that rely on end-to-end imitation learning from human demonstrations, where vision recognition and action planning are tightly integrated, our method divides these functions into distinct modules. This separation allows our agent to leverage the knowledge and reasoning power of the LLM, enabling it to handle unseen tasks without requiring additional training, as long as those tasks don’t introduce entirely new types of UI elements that the visual observer module fails to recognize.
5.2 WebShop
WebShop (Yao et al., 2022b) is a simulated e-commerce web environment designed to evaluate agents’ performance on tasks that more closely resemble real-world scenarios. To successfully complete WebShop tasks, the agent must understand textual instructions and requirements for the target product and navigate various types of web pages to locate it, out of a pool of 1.18 million real products scraped from amazon.com. The environment provides a total of 12,087 instructions, divided into 10,587 for training, 1,000 for development, and 500 for testing. Among them, we utilized only the 500 test instructions for evaluation purposes.
Models and datasets
The web browser for the WebShop environment can be resized freely, and for convenience, we used a fixed resolution of 1276 × 1153 pixels for both data collection and evaluation. Unlike MiniWoB++, the dataset used to train the visual observer for WebShop was generated entirely through an automated process. Given an instruction, a simple helper agent we created extracted a search term, performed a search, and clicked the first item in the results. Screen images were captured throughout this process, and annotation data was collected using Selenium for each image, resulting in a total of 300 annotated samples. The decision-making model, supported action types, and function-calling approach are entirely consistent with those employed in the MiniWoB++ experiments. For the evaluation of the CAAP agent, a single human demonstration is applied universally to all instructions.
Experiment result
Table 2: Comparison with other image-only methods for solving WebShop problems.
Method Modality Image datasize Demo datasize Task score
Human Image - - 82.1
Pix2Act Image 0 1K 46.7
CAAP Image 0.3K 1 62.3
There has been limited research on constructing agents that rely solely on image inputs, and Pix2Act is the only method whose performance has been validated on the WebShop benchmark. Table 2 presents a comparison between our approach and Pix2Act on WebShop. (Comparisons with other agent-based methods that rely on additional HTML input, on top of image input, are presented in Section 6 of Appendix B.2.) While Pix2Act employed 1,012 human demonstrations to train its agent model, our approach relied on just a single human demonstration, along with 300 samples collected through a fully automated process for training the visual observer module. Despite the significantly reduced manual data collection effort, our agent achieved a task score of 62.3, a meaningful improvement over Pix2Act’s score of 46.7, with a margin of 15.6. Although this result does not yet reach human-level performance, it demonstrates the potential of our approach to solve real-world computer tasks with minimal human effort in data collection.
5.3 Ablation studies
CoT inducing instructions for CAAP prompting
@@ 5.3 Ablation studies
Figure 4 presents experiment results comparing the effects of human demonstrations and CoT inducing instructions in our CAAP prompting method. It is shown that the use of CoT inducing instructions results in a meaningful improvement in performance. Removing CoT inducing instructions from the prompt results in a 2.3% drop compared to the baseline, and a 3.8% drop when the human demo is absent. When only rationales are excluded from human demonstrations, there is a reduction in task-solving SR by 3.5% relative to the fully-prompted baseline. Excluding the whole human demonstrations leads to a decrease of 9.5%. Additionally, the removal of both CoT-inducing instructions and human demonstrations results in an decrease in SR by as much as 13.3%, demonstrating that effective ICL provided by our CAAP prompting scheme significantly enhances the decision-making capacity of the LLM model in use. The per-task result are depicted in Table 10 of Appendix J.1.
Comparison of the CAAP and the RCI prompting
Table 3: Performance comparison of the CAAP and RCI promptings. The RCI agent with CAAP is configured by replacing the prompting method of the original RCI agent with the CAAP approach.
RCI agent with CAAP RCI agent (original)
Average SR 0.840 0.769
Since our main experiments combine GUI-based perception and execution with the CAAP prompting technique, we conducted additional experiments to isolate the effect of CAAP prompting as part of an ablation study. Without modifying the input/output channels, we tested how the performance of the RCI agent (Kim et al., 2023) improves when its prompting mechanism is replaced by our CAAP prompting. We first executed the original RCI code from the authors‘ official GitHub repository without modifications, testing only the 47 tasks listed in the repository by 50 times and taking the average. ***The success rate of the RCI agent we observed deviates from the results reported in the literature, despite our diligent efforts to replicate the experimental setup exactly. Upon inquiry, the RCI authors indicated that they no longer had access to the complete experimental logs used in their publication. Then, we replaced the prompting method with CAAP and reran the tests. To ensure a fair comparison, we used identical seeds for task instances, the same demo set provided by the repository, and the same LLM for queries.
As described in Table 3, the results demonstrate that applying CAAP prompting to the RCI agent significantly increased the average task-solving SR by 7.1%, from 76.9% to 84.0%. (Task-level results can be seen in Table 11 of Appendix J.2.) The prompting methods of RCI and CAAP exhibit two key differences. First, the RCI approach tends to modify even correct action plans during the iterative critique and revision process. In contrast, the CAAP method, by focusing on thorough context review rather than critique, is able to avoid this issue when formulating action plans. Second, RCI executes actions sequentially based on a fixed initial plan, failing to adapt to changes in the environment that occur during execution. Conversely, the CAAP method is designed to account for state changes resulting from prior actions. This experiment underscores the effectiveness of our CAAP prompting, confirming its potential to enhance the performance of LLMs in computer tasks.
@@ 6 Conclusion
In this paper, we introduce a modular agent designed to mimic how humans solve tasks on computers. Specifically, we aim to address the limitations of existing models that rely on imitation learning with large quantities of human demonstrations or non-universal input sources such as HTML and DOM. Our agent operates solely on visual inputs and significantly reduces the need for human demonstrations by leveraging the common knowledge and reasoning capabilities built into LLMs. Additionally, we introduce the CAAP prompting mechanism to enhance the decision-making capabilities of LLM-based agents on computer tasks. Evaluations on the MiniWoB++ and WebShop benchmarks demonstrate that our agent outperforms previous methods while using significantly less data. By adopting a modular approach that separates visual perception, action planning, and action execution, our method allows for easy adaptation to new tasks or model upgrades. Additionally, our agent is designed to operate beyond web environments and across inter-application settings, offering broader applicability and significant advantages in terms of generality.
@@
Appendix A MiniWoB++ Experiment Details
Models for visual observer
We utilize Ultralytics YOLOv8.0.143 (ult, ), employing the default hyperparameters with the exception of ’box’, ’cls’, ’dfl’, and ’max_patches’ set to 0.9, 0.05, 0.05, and 512, respectively. For Pix2Struct, we use the base model provided at https://huggingface.co/google/pix2struct-base, without any modifications to the default configuration.